# Emergence of spin-phonon coupling in Gd-doped Y2CoMnO6 double perovskite oxide: a combined experimental and ab-initio study

### v2

@@ 3 Results and discussion
Our calculations indicate that with the introduction of disorder in the system, the shift between prominent modes further increases approximately by 3%. To investigate the cause behind the enhanced spin-phonon coupling in the case of disordered YGCMO, we performed a detailed structural analysis of the sample in the presence/absence of disorder. It was observed that in the ordered sample the Co-O-Mn bond angle was enhanced by 0.46∘ as the magnetic configuration was changed from FM to C-type. However, the same was found to be reduced by 1.88∘ for the disordered sample (refer to Table 2). Analysis of the bond lengths of the CoO6 and MnO6 octahedra showed that in ordered YGCMO the average Co-O bond length is found to increase by 0.0117 Å and the average Mn-O bond length was found to decrease by 0.0044 Å (refer to Table 3) as the magnetic configuration was changed from FM to C-type AFM. On the contrary, for similar magnetic transformations, both the average Co-O and Mn-O bond lengths were found to get enhanced and reduced by 0.0021 Å and 0.0164 Å , respectively (refer to Table 3). However, there is in general an increase in the distortion of the CoO6 and MnO6 octahedra with change in magnetic ordering, which is found to increase with the introduction of anti-site disorder. It has been observed previously [61] that the relation between the frequencies of the RAMAN modes () and the bond length (l) is . Therefore, the modulation in bond length of a specific atom engaged with a particular vibration can lead to shifts in mode positions. This is also evident in our first-principles calculations which show that in the presence of anti-site disorder, the magnetism induced distortion of the lattice is more which eventually gets reflected in the phonon spectra.
@@ 4 Conclusion
In summary, temperature-dependent Raman study confirms that the sample holds its symmetry throughout the investigating temperature range. Anomalous softening of Ag stretching mode implies the appearance of SPC due to the presence of both FM and AFM interactions. Again, the nature of phonon linewidth curves and the insulating nature of the material discards the effect of magnetostriction on the observed softening of Ag mode. On the other hand, shift in mode position from the anharmonic approach exhibits an unconventional behavior towards the square of the magnetization. This is because of competing FM and AFM magnetic ordering, which is driven by ASD. The calculated value of the SPC strength is found to be 0.29 cm-1. Our DFT based first-principles calculations indicate the presence of SPC in both ordered and disordered YCMO. Interestingly, SPC is found to be enhanced in the presence of anti-site disorder at the Co/Mn sites. Microscopic analysis indicates the intricate relationship between magnetism, state of orderliness at the crystallographic sites and the crystal structure of the YGCMO sample. Such materials are known to hold great promise in the field of spintronics.

### v3

@@ 3 Results and discussion
In order to understand the nature of the influence of Gd3+ substitution at Y3+ sites on the Raman spectra, We calculated the total phonon density of states (TPDOS) for disordered YCMO and YGCMO. It is to be noted that analysis of the partial phonon density of states (PDOS) of YGCMO (as presented in Section S6(f) of the Supplementary Information) indicates that the translational and the librational modes related to the Gd3+ ion appear in the lower wave number region (as Gd3+ is heavier than Y3+ by 76.88% and wavenumber= (=mass)). A comparison of our calculated total phonon density of states for YCMO and YGCMO as presented in Fig. 8(a) shows that the overall TPDOS gets shifted to lower frequencies with Gd doping. This happens because of change in bond length and bond angle with Gd doping (refer to Table 2). Thus, we conclude that A-site doping with a cation of a larger ionic radius shifts the phonon modes to lower frequencies.
| Bond angle | Bond length | ||||
| Co-O-Mn (∘) | Co-O-Co (∘) | Mn-O-Mn (∘) | Co-O (Å) | Mn-O (Å) | |
| disordered YCMO | 146.34 | 144.31 | 142.92 | 2.0241 | 1.9543 |
| disordered YGCMO | 146.66 | 147.28 | 145.16 | 2.0426 | 1.9731 |
With 50% of the Y being replaced by Gd, one can very well appreciate that it will give rise to various different configurations depending upon how the Gd is arranged in the YCMO lattice. Hence, we have investigated the influence of Gd ordering on the Raman spectra of YGCMO. Apart from Gd-Y ordering at A site, we also encounter Co-Mn ordering at the B site of the double perovskite. Our calculations reveal that any alteration in Gd ordering modifies the total phonon density of states (TPDOS) or Raman spectra (refer to S8 with in the SI).
We also calculated the PDOS of ordered and disorderd YGCMO to unveil the effect of antisite disorder on Raman spectrum. As illustrated in Fig. 8 (b), the phonon density of states (PDOS) for ordered and disordered YGCMO shows notable differences across the frequency spectrum. Antisite disorder at the B-site introduces Co-O-Co and Mn-O-Mn bonds, which increase lattice strain. This disorder modifies bond lengths and angles (refer to Table 3 and Table 4) (as wavenumber (=bond length)), thereby altering the PDOS. These structural changes directly influence the Raman spectrum.
@@
Our calculations indicate that with the introduction of disorder in the system, the shift between prominent modes further increases approximately by 3%, whereas it is found that the shift has been increased with an increase in Gd concentration. Specifically, for Gd2CoMnO6, the shift rises by 7% as shown in S9 (b) of the SI, indicating an enhanced spin-phonon coupling (SPC). To investigate the cause behind the enhanced spin-phonon coupling in the case of disordered YGCMO, we performed a detailed structural analysis of the sample in the presence/absence of disorder. It was observed that in the ordered sample the Co-O-Mn bond angle was enhanced by 0.46∘ as the magnetic configuration was changed from FM to C-type. However, the same was found to be reduced by 1.88∘ for the disordered sample (refer to Table 3). Analysis of the bond lengths of the CoO6 and MnO6 octahedra showed that in ordered YGCMO the average Co-O bond length is found to increase by 0.0117 Å and the average Mn-O bond length was found to decrease by 0.0044 Å (refer to Table 4) as the magnetic configuration was changed from FM to C-type AFM. On the contrary, for similar magnetic transformations, both the average Co-O and Mn-O bond lengths were found to get enhanced and reduced by 0.0021 Å and 0.0164 Å , respectively (refer to Table 4). However, there is in general an increase in the distortion of the CoO6 and MnO6 octahedra with change in magnetic ordering, which is found to increase with the introduction of anti-site disorder. It has been observed previously [66] that the relation between the frequencies of the Raman modes () and the bond length () is . Therefore, the modulation in bond length of a specific atom engaged with a particular vibration can lead to shifts in mode positions. This is also evident in our first-principles calculations, which show that in the presence of anti-site disorder, the magnetism induced distortion of the lattice is more, which eventually gets reflected in the phonon spectra.
@@ 4 Conclusion
In summary, temperature-dependent Raman study confirms that the sample holds its symmetry throughout the investigated temperature range. The anomalous softening of Ag stretching mode implies an emergence of SPC from the presence of both FM and AFM interactions. Again, the nature of phonon linewidth curves and the half-metallic nature of the material repudiate any effect of magnetostriction on the observed softening of Ag modes. On the other hand, shift in the Ag mode position from the anharmonic approach exhibits an unconventional behavior with the square of the magnetization. This is because of competing FM and AFM magnetic ordering, which is driven by ASD. We have found from DFT calculations that the Gd substitution at A-site induces a shift of phonon modes toward the lower frequency region. Additionally, antisite disorder (ASD) and Gd ordering at the A-site significantly modify the Raman spectra. Our calculations also indicate the presence of SPC in both ordered and disordered YGCMO. The calculated value of the SPC strength is 0.29 cm-1. Notably, Gd doping at the A-site and ASD at the B-site significantly enhance SPC in the system. Microscopic analysis indicates an intricate relationship between magnetism, state of orderliness at the crystallographic sites, and the crystal structure of the YGCMO sample. Such materials are known to hold promise in the field of spintronics.